OpenStudy (anonymous):

what is the period of f(x)=cot pix/8

OpenStudy (anonymous):

period is 8

OpenStudy (anonymous):

in general if you have \[\cot(bx)\] period is \[\frac{\pi}{b}\] so in this case you get \[\frac{\pi}{\frac{\pi}{8}}=8\]

OpenStudy (anonymous):

AAH thank you so much for your help!